OpenStudy (anonymous):

What is the solution set of |x + 3| - 2 = -5? A- -6 or 0 B- No solution C- -5 or 5 D- -3 or 3

jigglypuff314 (jigglypuff314):

the answer is A because when |x+3|-2=-5 +2 +2 |x+3| =-3 then the absolute value factors out to -(x+3)=-3 or (x+3)=-3 - x -3 =-3 x+3 =-3 +3 +3 -3 -3 -x = 0 x = -6 x = 0 so x = -6 or 0

OpenStudy (anonymous):

Thank You!! :D